OpenStudy (anonymous):

250 square foot section with a fence that costs $1.50 per yard. What dimensions would you use to minimize the cost? In your own words, please post a response

OpenStudy (anonymous):

Lets make the dimensions of this area x for length, and y for width. Since the area is 250 square feet, we have: \[xy = 250 \Rightarrow y = \frac{250}{x}\] At the same time, the perimeter is going to be: \[2(x+y) \Rightarrow 2(x+\frac{260}{x})\] To minimize cost, i want to minimize the perimeter, so im going to take the derivative and set it equal to zero: \[P = 2x+\frac{500}{x} \Rightarrow 0 = 2-\frac{500}{x^2}\] Now we just solve for x: \[0 = 2-\frac{500}{x^2} \Rightarrow 0 = 2x^2-500 \Rightarrow 2x^2 = 500 \Rightarrow x^2 = 250 \Rightarrow x = 5\sqrt{10}\] The last part is figuring out what y is: \[y = \frac{250}{x} = \frac{250}{5\sqrt{10}} = \frac{50}{\sqrt{10}} = \frac{50\sqrt{10}}{10} = 5\sqrt{10}\]

OpenStudy (anonymous):

Joe thank you